OpenStudy (anonymous):

simplify (7 – c)(–1)

OpenStudy (anonymous):

-7+c..you have to note that -multiplied by - is + but - and + is -

OpenStudy (anonymous):

(7-c)(-1) =-7+c or =c-7 use the distributive property